# Privacy Reasoning in Ambiguous Contexts

### v2

@@ 6 Results
Many of the reasoning-guided codes we develop (see Table 2) apply to only a small subset of examples, but we observe in Figure 4 that expanding context based on a single code significantly improves performance across all examples. To understand how these codes generalize, we split the 120 coded examples in PrivacyLens+ into two groups, those labeled with a particular code and those without. Figure 8 shows the performance of those two groups before and after expansion using each of the 9 codes. Perhaps surprisingly, the performance gains (measured by the large percentage of wrong -> right group) on examples labeled with a particular code are not significantly higher than the performance gains on those without (i.e., clarification in one code generalizes in terms of helping in examples where the LLM-generated reasoning stated other assumptions). This shows that on a per-example basis, expanding based on the code mentioned in LLM-generated reasoning does not outperform choosing any other codes from the codebook. This is likely due to LLM-provided reasoning not reflecting the underlying reasoning that led to the specific judgment, as also reported in prior work (chen2025reasoning). Additionally, many labeled examples may have more than one code assigned to them, making it harder to dissect the effect of one particular code. Nevertheless, model-generated reasoning in aggregate reflects the underlying considerations in the LLM reasoning process, as supported by the observation that for all codes, the expansions correct most of the previously incorrectly judged examples (i.e., 27% - 70% examples are wrong -> right). This suggests that while aggregate reasoning provides a powerful signal for identifying general areas of ambiguity, the specific rationale provided for any single decision may not be a reliable guide to the model’s underlying inference process, highlighting a critical challenge in LLM interpretability. A qualitative analysis of the expansions is performed in Appendix B.3.
@@ 7 Discussion and Limitations
In Section 5, we consider three strategies to disambiguate the context present in the scenarios in our datasets: providing additional label-independent context, label-dependent contexts based on the definition of fields, and label-dependent context designed to address assumptions identified in the model’s reasoning (Section 5.2). We observe significant accuracy gains and reductions in prompt sensitivity, which suggest that research on methods to disambiguate context is a crucial area for future work and that such methods can complement model tuning. Across these experiments, we see the largest improvements in performance—and reductions in prompt sensitivity—when addressing assumptions made by the model. In particular, model reasoning to disambiguate context outperforms the disambiguation based on contextual integrity parameters, which the theory identifies as the key determinants of privacy decision-making nissenbaum2004privacy. This suggests that the guidance derived from model reasoning is helpful for creating more targeted questions or a more fine-grained taxonomy of the salient information that is critical for privacy decision-making.
@@ B.2.4 Performance generalizes across prompt variants
To further show the benefits of disambiguation, we performed privacy judgments using 4 additional fine-grained prompt variants, inspired by prior work (shvartzshnaider2024position), that paraphrase the original neutral prompt. The 4 variants, which are provided verbatim for both PrivacyLens+ and ConfAIde+ in Appendix E, are:
@@ C.4 Reasoning-guided expansions
To understand model reasoning and identify implicit assumptions, two authors analyzed 40 outputs for PrivacyLens+ (10 per category: TPs, TNs, FPs, FNs) and 21 outputs for ConfAIde+ (7 each for TPs, TNs, FNs; FPs are excluded due to lack of samples). They examine the prompts and reasoning for those examples – blind to ground truth labels and LLM judgments – to identify assumptions and develop 9 codes shown in Table 2. Subsequently, using this established codebook, the same two authors independently coded a larger, stratified random sample of 120 PrivacyLens+ examples (30 per category) and 50 ConfAIde+ examples (15 TPs, TNs, FNs, and 5 FPs). Disagreements were resolved through discussion to achieve consensus cascio2019team. Table 2 presents the resulting code occurrences across datasets.
@@
After identifying potential assumptions made by the LLM, we use them to add clarifying context. The goal of this expansion is to help the LLMs arrive at the right answer by reducing implicit reasoning assumptions. We denote the label-dependent contexts we add to the $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$ examples as, respectively $\mathcal{N^{-}_{C}}$ and $\mathcal{N^{+}_{C}}$. To generate an $\mathcal{N^{-}_{C}}$ (or $\mathcal{N^{+}_{C}}$) for a specific identified code, the LLM is instructed to first select the most suitable field for expanding the context for the code and its definition, then expand the selected field strictly following the code and its definition to make the example more inappropriate (or appropriate, respectively). Similarly to Shao et al. salt-nlp2024privacylens, we explicitly ask the LLM to add descriptive contexts and avoid using evaluative words like “sensitive” and “non-sensitive” that are indicative of appropriateness that would introduce reasoning shortcuts in the expansions. This process is repeated for all $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$ examples in PrivacyLens+, adding an additional $9\times 493$ = 4,437 pairs of $\mathcal{S^{-}_{N^{-}_{C}}}$ and $\mathcal{S^{+}_{N^{+}_{C}}}$ examples to the dataset. For ConfAIde+, the LLM is additionally instructed to incorporate generated $\mathcal{N^{-}_{C}}$ and $\mathcal{N^{+}_{C}}$ into the original $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$ examples, creating an additional $9\times 270$ = 2,430 pairs of $\mathcal{S^{-}_{N^{-}_{C}}}$ and $\mathcal{S^{+}_{N^{+}_{C}}}$ examples. The generation prompts can be found in Appendix D.1.3 and D.2.4.
@@ Appendix D Datasets and dataset expansions
Several benchmarks for probing-style LLM privacy assessments mireshghallah2023confaide; wang2023decodingtrust; sun2024trustllm; zhang2024multitrust focus on preventing leakage of sensitive data. This means, however, that they contain only inappropriate examples; i.e., for the data sharing scenarios and data-probing questions in these datasets, the data should not be shared. This enables testing the extent to which models are willing to leak sensitive data or refuse to answer, but does not test the utility of a model—or agent—in contexts in which sensitive data is perfectly acceptable to share. Lacking a measure of utility in addition to privacy leakage does not provide a complete picture of the performance of various agents: a model that always answers ’no’ would perfectly protect privacy, yet its practical utility at revealing contextually appropriate information would be low. While more recent benchmarks salt-nlp2024privacylens focus on capturing the helpfulness of actions and trajectories of agents, they do not focus on the probing privacy awareness – a task that requires reasoning capabilities beyond privacy mireshghallah2023confaide.
@@
ConfAIde contains four tiers for privacy assessments of LLMs, ranging from non-contextual, to task-oriented agent scenarios. Among these, tier 3 focuses on probing-style questions to determine various facets of secret keeping and theory-of-mind in LLMs. This benchmark, which contains scenarios involving three actors, the first (subject agent) disclosing a sensitive personal detail with one of the second (aware agent), and the second having the opportunity to disclose this detail with the third actor (oblivious agent) for an incentive (reveal reason). In addition, each example also contains the relationship between subject agent and the other agents, (aware-agent relationship, oblivious-agent relationship). We refer to this data structure as the seed. Because these fields do not contain sufficient contextual information, they are expanded into scenarios (short stories based on a template) using an LLM. The privacy judgment involves probing the model on whether it is okay to share the secret given the scenarios. The complete version of the judgment prompt is in Prompt 21 from Appendix E. The answer to these questions should always be ’no’ and any deviation from this answer represents a privacy leakage through the evaluated models. The benchmark contains 270 examples of negative scenarios $\mathcal{S^{-}}$, and has been used extensively by prior work shvartzshnaider2024llm; sun2024trustllm; zhang2024multitrust.
@@ D.1 PrivacyLens+
The privacy-sensitive seed in PrivacyLens are constructed based on privacy norms collected from legal regulations, academic privacy literature and findings from their own Prolific-based crowdsourcing. Developed based on the contextual integrity theory (nissenbaum2004privacy; nissenbaum2009privacy), these seeds distill potential privacy violations into structured components, effectively removing narrative complexities, making them an effective data structure for evaluating how well LLMs answer probing-style questions about information sharing. An example seed is shown in Example 1. All examples in PrivacyLens+ are developed based on these seed.

### v3

@@ 6 Results
Many of the reasoning-guided codes we develop (see Table 2) apply to only a small subset of examples, but we observe in Figure 4 that expanding context based on a single code significantly improves performance across all examples. To understand how these codes generalize, we split the 120 coded examples in PrivacyLens+ into two groups, those labeled with a particular code and those without. Figure 8 shows the performance of those two groups before and after expansion using each of the 9 codes. Perhaps surprisingly, the performance gains (measured by the large percentage of wrong -> right group) on examples labeled with a particular code are not significantly higher than the performance gains on those without (i.e., clarification in one code generalizes in terms of helping in examples where the LLM-generated reasoning stated other assumptions). This shows that on a per-example basis, expanding based on the code mentioned in LLM-generated reasoning does not outperform choosing any other codes from the codebook. This is likely due to LLM-provided reasoning not reflecting the underlying reasoning that led to the specific judgment, as also reported in prior work (Chen et al., 2025). Additionally, many labeled examples may have more than one code assigned to them, making it harder to dissect the effect of one particular code. Nevertheless, model-generated reasoning in aggregate reflects the underlying considerations in the LLM reasoning process, as supported by the observation that for all codes, the expansions correct most of the previously incorrectly judged examples (i.e., 27% - 70% examples are wrong -> right). This suggests that while aggregate reasoning provides a powerful signal for identifying general areas of ambiguity, the specific rationale provided for any single decision may not be a reliable guide to the model’s underlying inference process, highlighting a critical challenge in LLM interpretability. A qualitative analysis of the expansions is performed in Appendix B.3.
@@ 7 Discussion and Limitations
In Section 5, we consider three strategies to disambiguate the context present in the scenarios in our datasets: providing additional label-independent context, label-dependent contexts based on the definition of fields, and label-dependent context designed to address assumptions identified in the model’s reasoning (Section 5.2). We observe significant accuracy gains and reductions in prompt sensitivity, which suggest that research on methods to disambiguate context is a crucial area for future work and that such methods can complement model tuning. Across these experiments, we see the largest improvements in performance—and reductions in prompt sensitivity—when addressing assumptions made by the model. In particular, model reasoning to disambiguate context outperforms the disambiguation based on contextual integrity parameters, which the theory identifies as the key determinants of privacy decision-making Nissenbaum (2004). This suggests that the guidance derived from model reasoning is helpful for creating more targeted questions or a more fine-grained taxonomy of the salient information that is critical for privacy decision-making.
@@ B.2.4 Performance generalizes across prompt variants
To further show the benefits of disambiguation, we performed privacy judgments using 4 additional fine-grained prompt variants, inspired by prior work (Shvartzshnaider and Duddu, 2024), that paraphrase the original neutral prompt. The 4 variants, which are provided verbatim for both PrivacyLens+ and ConfAIde+ in Appendix E, are:
@@ C.4 Reasoning-guided expansions
To understand model reasoning and identify implicit assumptions, two authors analyzed 40 outputs for PrivacyLens+ (10 per category: TPs, TNs, FPs, FNs) and 21 outputs for ConfAIde+ (7 each for TPs, TNs, FNs; FPs are excluded due to lack of samples). They examine the prompts and reasoning for those examples – blind to ground truth labels and LLM judgments – to identify assumptions and develop 9 codes shown in Table 2. Subsequently, using this established codebook, the same two authors independently coded a larger, stratified random sample of 120 PrivacyLens+ examples (30 per category) and 50 ConfAIde+ examples (15 TPs, TNs, FNs, and 5 FPs). Disagreements were resolved through discussion to achieve consensus Cascio et al. (2019). Table 2 presents the resulting code occurrences across datasets.
@@
After identifying potential assumptions made by the LLM, we use them to add clarifying context. The goal of this expansion is to help the LLMs arrive at the right answer by reducing implicit reasoning assumptions. We denote the label-dependent contexts we add to the $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$ examples as, respectively $\mathcal{N^{-}_{C}}$ and $\mathcal{N^{+}_{C}}$. To generate an $\mathcal{N^{-}_{C}}$ (or $\mathcal{N^{+}_{C}}$) for a specific identified code, the LLM is instructed to first select the most suitable field for expanding the context for the code and its definition, then expand the selected field strictly following the code and its definition to make the example more inappropriate (or appropriate, respectively). Similarly to Shao et al. Shao et al. (2024), we explicitly ask the LLM to add descriptive contexts and avoid using evaluative words like “sensitive” and “non-sensitive” that are indicative of appropriateness that would introduce reasoning shortcuts in the expansions. This process is repeated for all $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$ examples in PrivacyLens+, adding an additional $9\times 493$ = 4,437 pairs of $\mathcal{S^{-}_{N^{-}_{C}}}$ and $\mathcal{S^{+}_{N^{+}_{C}}}$ examples to the dataset. For ConfAIde+, the LLM is additionally instructed to incorporate generated $\mathcal{N^{-}_{C}}$ and $\mathcal{N^{+}_{C}}$ into the original $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$ examples, creating an additional $9\times 270$ = 2,430 pairs of $\mathcal{S^{-}_{N^{-}_{C}}}$ and $\mathcal{S^{+}_{N^{+}_{C}}}$ examples. The generation prompts can be found in Appendix D.1.3 and D.2.4.
@@ Appendix D Datasets and dataset expansions
Several benchmarks for probing-style LLM privacy assessments Mireshghallah et al. (2024); Wang et al. (2023); Sun et al. (2024); Zhang et al. (2024) focus on preventing leakage of sensitive data. This means, however, that they contain only inappropriate examples; i.e., for the data sharing scenarios and data-probing questions in these datasets, the data should not be shared. This enables testing the extent to which models are willing to leak sensitive data or refuse to answer, but does not test the utility of a model—or agent—in contexts in which sensitive data is perfectly acceptable to share. Lacking a measure of utility in addition to privacy leakage does not provide a complete picture of the performance of various agents: a model that always answers ’no’ would perfectly protect privacy, yet its practical utility at revealing contextually appropriate information would be low. While more recent benchmarks Shao et al. (2024) focus on capturing the helpfulness of actions and trajectories of agents, they do not focus on the probing privacy awareness – a task that requires reasoning capabilities beyond privacy Mireshghallah et al. (2024).
@@
ConfAIde contains four tiers for privacy assessments of LLMs, ranging from non-contextual, to task-oriented agent scenarios. Among these, tier 3 focuses on probing-style questions to determine various facets of secret keeping and theory-of-mind in LLMs. This benchmark, which contains scenarios involving three actors, the first (subject agent) disclosing a sensitive personal detail with one of the second (aware agent), and the second having the opportunity to disclose this detail with the third actor (oblivious agent) for an incentive (reveal reason). In addition, each example also contains the relationship between subject agent and the other agents, (aware-agent relationship, oblivious-agent relationship). We refer to this data structure as the seed. Because these fields do not contain sufficient contextual information, they are expanded into scenarios (short stories based on a template) using an LLM. The privacy judgment involves probing the model on whether it is okay to share the secret given the scenarios. The complete version of the judgment prompt is in Prompt 21 from Appendix E. The answer to these questions should always be ’no’ and any deviation from this answer represents a privacy leakage through the evaluated models. The benchmark contains 270 examples of negative scenarios $\mathcal{S^{-}}$, and has been used extensively by prior work Shvartzshnaider et al. (2024); Sun et al. (2024); Zhang et al. (2024).
@@ D.1 PrivacyLens+
The privacy-sensitive seed in PrivacyLens are constructed based on privacy norms collected from legal regulations, academic privacy literature and findings from their own Prolific-based crowdsourcing. Developed based on the contextual integrity theory (Nissenbaum, 2004, 2009), these seeds distill potential privacy violations into structured components, effectively removing narrative complexities, making them an effective data structure for evaluating how well LLMs answer probing-style questions about information sharing. An example seed is shown in Example 1. All examples in PrivacyLens+ are developed based on these seed.